A ball of mass 0.1kg is thrown vertically upwards with an initial velocityof 80 m/s. calculate the pontential energy (i) half way up (ii) at its maximum height. what is its kinetic energy as it leaves the ground​.​
1
Expert's answer
2020-10-12T07:46:33-0400

The kinetic energy as it leaves the ground​:


"K=0.5mv^2=0.5(0.1)(80)^2=320\ J"

i)


"E=0.5K=0.5(320)=160\ J"

ii)


"E=K=320\ J"